HUBBLE SPACE TELESCOPE DAILY REPORT #2541 PERIOD COVERED: 0000Z (UTC) 01/17/00 - 0000Z (UTC) 01/18/00 Daily Status Report as of 018/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8411 (STIS CCD Spectroscopic Flats C-8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed Two Sets of FGS 8522 (Adjusting FGS2r's AMA for Optimal Interferometric Performance) The FGS2R was used to obtain S-curves at selected locations within the FGS2R FOV. These data will be will be used to adjust the AMA so as to align the HST pupil onto the interferometer's Koesters prisms. The final AMA position will be set to produce an acceptable set of s-curves across the FGS2r FOV. The proposal completed with no reported problems. 1.3 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed Three Sets of WF/PC-2 8491 (Transition to Hold, Decon and Cool Down Procedure) This program covers the transition from Protect Safe Mode to Hold Mode, then performs a Protect Decontamination. This is followed by a series of UV monitors and focus checks at -43C, then a special decon to take the instrument to -88C. The program completed nominally. 1.5 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed STIS/CCD 8412 (STIS CCD Imaging Flats C8) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field stability of the unit over a monthly period. A bias calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/MA1/MA2 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.8 Completed STIS/CCD/MA1 8107 (What is the Nature of the Cold Medium in Cooling Flow Clusters?) The Space Telescope Imaging Spectrograph (CCD/MA1) was used to observe X-ray spectra of cooling flow clusters of galaxies in an attempt to discover the presence of large columns of "cold'' absorbing material. The proposal completed with no reported problems. 1.9 Completed STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.10 Completed STIS/CCD 8511 (Full-field Sensitivity and Geometry) The Space Telescope Imaging Telescope (CCD) was used to measure the geometric distortion for each STIS detector, using the same field as previously used for measuring the MAMA geometric distortion in cycle 7. These exposures will also serve to measure the image location for each of the MAMA imaging modes, with respect to the CCD frame, In order to fix the detector coordinates for the STIS MAMA imaging modes with respect to the CCD coordinates, and hence to the FGS system. The proposal completed nominally. 1.11 Completed FGS 8334 (HD 98800: An Opportunity to Measure True Masses for Low-Mass PMS Stars) The Fine Guidance Sensors were used to make astrometric observations of the low mass pre-main sequence (PMS) star HD98800 in order to measure its true mass. HD 98800 became interesting when IRAS found it to have a large infrared excess, indicating a substantial dust disk. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-22 8498 (Internal Monitor - Post-TEC Turn On) The WF/PC-2 was used to obtain routine internal observations in order to monitor basic instrument health, and optical alignment (K-spots, INTFLATS, VISFLATS, UVFLATS, bias, Darks). The observations completed nominally. 1.13 Completed STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ1651+4+5 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 3 Successful: 3 2.2 FHST Updates: Scheduled: 23 Successful: 23 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) activities dominate the observations currently being performed. Using ROP SR-1A, the SSR EDAC error counter was reset once. Per HSTAR 7475, on five occasions, beginning at 017/122358Z and lasting for approximately nine minutes, Coarse Sun Sensor (CSS) #2 position parameter (GAPXASS2) flagged out-of-limits low with a value of -5.04001. There was no CCS #2 sun presence at the time. With ROP DF-07A as a guide, the ephemeris tables were uplinked at 018/0007Z. The acquisition at 018/030811Z defaulted to fine lock backup on FGS-1 only when the scan step limit was exceeded. HSTAR 7476 was written. STIS/CCD/MA1 proposal 8511, occurring during the period of the next report, may have been affected. The STIS EMC experienced a re-try at 108/034146Z. Per ROP NS-12, the STIS flight software error counter was reset at 018/0501Z. At 018/060928Z, various coarse sun sensor parameters (GAPCASS1, GAPXASS2, GAPCBSS1, and GAPXBSS2) flagged out-of-limits with values of 5.04 and -5.04. Limits are 5.0 and -5.0. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW